The angle a kite string makes with the ground is 24º. If 80 m of string has been let out, how high is the kite above the ground?
The angles of elevation from a point A and a point B to the top of a mountain 780m high are shown. Point A is 67 degrees and point B is 48 degrees. How long would a tunnel be through the mountain, connecting A and B?

Answer by mananth(16949)

The kite is vertically above the ground. height ?
The length of sting = hypotenuse= 80m
The angle of elevation = 24 deg.
Sin theta = opposite side/ hypotenuse
sin 24 = opposite side/80
height of kite = 80 * sin 24
use calc to find the height
.....
Point A elevation is 67 deg
point B elevation = 48 deg.
height of mountain = 780 m
A to mountain foot distance = 780/Tan 67 .............. (1)
B to mountain foot distance = 780/tan 48 ..................(2)
distance AB = (2)-(1)
780/1.11 -780/2.35
about 370.8 m is the length of the tunnel.